A wooden log of 0.8 m diameter and 6 m length is floating in a river water. Find the depth of wooden log in water when the sp. gr. Of the wooden log is 0.7.


1
Expert's answer
2021-11-06T01:35:25-0400
"mg=\rho_wgV_\text{submerged},\\\space\\
V_\text{submerged}=\frac{m}{\rho_w}=\frac{SG·\rho_w·V}{\rho_w},\\\space\\
V_\text{submerged}=SG·V=SG·\frac{\pi d^2}{4}L."

The volume above the water surface is


"\Delta V=V-V_\text{submerged}=(1-SG)·\frac{\pi d^2}{4}L."

Divide by length to find the area of the segment above the water:


"A=(1-SG)·\frac{\pi d^2}{4}."


The area of the segment can be expressed in terms of areas of a triangle and sector:


"A=A_s-A_t=\pi\theta-x·(r-h),\\
\theta=2\arctan\frac xh.\\\space\\
A=2\pi\arctan\frac xh-x·(r-h).\\\space\\
x^2=r^2-(r-h)^2=2rh-h^2.\\\space\\
h=0.26\text{ m}."

Therefore, the depth is


"D=d-h=0.54\text{ m}."